### v1

# A greedy heuristic for graph burning

Abstract
Given a graph $G$, the optimization version of the graph burning problem seeks for a sequence of vertices, $(u_{1},u_{2},...,u_{k})\in V(G)^{k}$, with minimum $k$ and such that every $v\in V(G)$ has distance at most $k-i$ to some vertex $u_{i}$. The length $k$ of the optimal solution is known as the burning number and is denoted by $b(G)$, an invariant that helps quantify the graph’s vulnerability to contagion. This paper explores the advantages and limitations of an $\mathcal{O}(mn+kn^{2})$ deterministic greedy heuristic for this problem, where $n$ is the graph’s order, $m$ is the graph’s size, and $k$ is a guess on $b(G)$. This heuristic is based on the relationship between the graph burning problem and the clustered maximum coverage problem, and despite having limitations on paths and cycles, it found most of the optimal and best-known solutions of benchmark and synthetic graphs with up to 102400 vertices.
Keywords graph burning  $\cdot$ clustered maximum coverage  $\cdot$ approximation algorithm  $\cdot$ greedy heuristic  $\cdot$ social contagion
1 Introduction
The optimization version of the graph burning problem (GBP) is an NP-hard minimization problem that helps quantify a graph’s vulnerability to contagion. It was introduced as a social contagion model [1] and a processor communication problem [2]. Nevertheless, it is an abstract problem with application in many other contexts, particularly those related to information or influence spreading. In detail, GBP receives a graph $G$ as input and seeks a burning sequence of minimum length, where a burning sequence is a list of vertices that burns all the graph’s vertices by the end of the burning process. The latter consists of burning the neighbors of the previously burned vertices and the current vertex in the sequence at discrete time steps. Figure 1 shows the burning process of the optimal burning sequence of a given input graph. The length of the optimal burning sequence of a graph $G$ is known as its burning number and is denoted by $b(G)$.
Figure 1: Burning process of the optimal burning sequence $(v_{4},v_{8},v_{1})$; notice that $b(G)=3$. At each discrete time step, the fire propagates to the neighbors of the previously burned vertices, and the next vertex in the sequence gets burned too. In the end, all vertices are burned.
GBP has an interesting relationship with other NP-hard problems, such as minimum dominating set (MDSP) [3], minimum set cover (MSCP) [4], vertex k-center (VKCP) [5, 6, 7, 8], and firefighter (FP) problems [9]. MDSP can be seen as the GBP with a parallel diffusive process; all vertices are burned in one discrete time step. VKCP reduces to a series of MDSPs and, by transitivity, to a series of MSCPs [5]. Besides, its approximation algorithms are similar to those for GBP [10, 11, 12]. Regarding FP, it aims to protect vertices given the same propagation model used by GBP [9]. This paper shows how GBP is also related to a clustered maximum coverage problem (CMCP), a generalization of the NP-hard maximum coverage problem (MCP) [13, 14]. The remaining part of the paper is organized as follows. Section 2 summarizes the related work, introduces an integer linear program for GBP, and lists the main definitions used in the following sections. Section 3 shows how GBP reduces to a series of CMCPs. Section 4 introduces a $(1/2)$-approximation algorithm for CMCP and a deterministic greedy heuristic for GBP based on this same approximation algorithm. The proposed heuristic is conceptually simple and easy to implement. Besides, it performed unexpectedly well on benchmark and synthetic graphs, as reported in Section 5.
2 Related work
The burning number conjecture (BNC) is one of the main open questions in the GBP literature. It states that the burning number of any connected graph $G$ is upper bounded by $\lceil n^{1/2}\rceil$ [15], where $n=|V(G)|$. The best-known bound for GBP over arbitrary connected graphs is $\lceil(4n/3)^{1/2}\rceil+1$ [16]. Much research has focused on complexity analysis and finding upper bounds for the burning number on specific graph families, including paths [17], path-forests [17, 11, 18], trees [17], grids [19], intervals [19], fences [20], theta graphs [21], spiders [18], caterpillars [22], graph products [23], generalized Petersen graphs [24], and Jahangir graphs [25]. Parameterized complexity has been studied too [26, 27]. Generalizations of GBP have also been proposed, such as the $w$-graph burning problem ($w$GBP), where $w$ vertices can be selected at each discrete time step [28]; the $r$-graph burning problem ($r$GBP), where a vertex gets burned only if it is adjacent to at least $r$ burned vertices [29]; and the independent cascade graph burning problem (ICGBP), where a vertex $v$ can burned by one of its neighbors $u$ only if the influence $u$ exerts over $v$ is larger than a given threshold [30]. Some heuristics, metaheuristics, approximation algorithms, and mathematical formulations have been proposed to find feasible solutions for GBP. Among the state-of-the-art heuristics and metaheuristics are a centrality-based genetic algorithm (CBAG) [31], a maximum eigenvector centrality heuristic (MECH) [32], a cutting corners heuristic (CCH) [32], a greedy algorithm with forward-looking search strategy (GFSSH) [32], a backbone-based greedy heuristic (BBGH) [33], an improved cutting corners heuristic (ICCH) [33], and a component-based recursive heuristic (CBRH) [33]. Through experimental results, these have proven effective in finding feasible solutions. Among the approximation algorithms are proposals with an approximation ratio of 3 [11] and $3-2/b(G)$ [12] for arbitrary graphs, of 2 for trees [11] and square grids [19], of 1.5 for disjoint paths [11], and polynomial time approximation schemes (PTAS) for linear forests [11] and graphs with small treewidth [34]. Regarding mathematical formulations, an integer linear program (ILP) and two constraint satisfaction problems (CSPs) have been reported [35]. To complement the formal statement of the problem, let us introduce the following ILP, which is equivalent to but simpler than the one previously reported in the literature [35].
min $\ \ \sum\limits_{j\in[1,U]}b_{j}$ (1)
s.t. $\ \ b_{j}\geq 1-b_{ij}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (2)
$\ \ b_{ij}\leq s_{ij}+\sum\limits_{v_{k}\in N[v_{i}]}b_{kj-1}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (3)
$\ \ \sum\limits_{v_{i}\in V}s_{ij}=w$ $\forall j\in[1,U]$ (4)
where $\ \ s_{ij}\ ,\ b_{ij}\ ,\ b_{j}\in\{0,1\}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (5)
$\ \ b_{i0}=0$ $\forall v_{i}\in V$ (6)
Given an input graph $G=(V,E)$, let $n=|V|$, $m=|E|$, and $V=\{v_{1},v_{2},...,v_{n}\}$. This ILP has $\mathcal{O}(Un)$ binary decision variables and $\mathcal{O}(Un)$ constraints, where $U$ is an upper bound on $b(G)$, i.e., $b(G)\leq U\leq n$. Variables $b_{ij}$ (resp., $s_{ij}$) codify which vertices $v_{i}$ are burned (resp., added to the sequence) at each discrete time step $j$. Specifically, $b_{ij}=1$ (resp., $s_{ij}=1$) if vertex $v_{i}$ is burned (resp., is in the sequence) at the discrete time step $j$; $b_{ij}=0$ (resp., $s_{ij}=0$) otherwise. The constant $w$ serves to generalize the program to $w$GBP. If $w=1$, this ILP corresponds to the original GBP. Constraints (2) set $b_{j}=1$ if some vertex is not burned at the discrete time step $j$; since the Objective function (1) is minimized, if all vertices are burned at the discrete time step $j$, then $b_{j}=0$. This way, the objective function seeks to burn all vertices as soon as possible, i.e., at the smallest index $j$. Constraints (3) guarantee that a vertex cannot be burned unless it is into the sequence, was already burned, or has some burned neighbor. Finally, Constraints (4) indicate that $w$ vertices must be added into the sequence at each discrete time step. The solution is codified into variables $s_{ij}$ and $b(G)$ equals the smallest index $j$ at which all vertices are burned; namely, $b(G)$ is the value of the objective function plus one. Before finishing this section, let us list the basic definitions used in the remaining part of the paper.
Basic definitions
Definition 1.
A graph $G=(V,E)$ is an ordered pair, where $V$ is the set of vertices and $E$ is the set of edges, a set of 2-element subsets of $V$.
Definition 2.
The distance between two vertices in a graph is the length of their shortest path.
Definition 3.
The open neighborhood $N(v)$ of a vertex $v$ is its set of neighbors in $G$.
Definition 4.
The closed neighborhood $N[v]$ of a vertex $v$ is $N(v)\cup\{v\}$.
Definition 5.
The $r^{th}$ power $G^{r}$ of a given graph $G=(V,E)$ is the graph that results from adding an edge between each pair of vertices in $V$ at a distance at most $r$.
Definition 6.
The $r^{th}$ open neighborhood $N_{r}(v)$ of a vertex $v$ is its set of neighbors in $G^{r}$.
Definition 7.
The $r^{th}$ closed neighborhood $N_{r}[v]$ of a vertex $v$ is $N_{r}(v)\cup\{v\}$.
Definition 8.
A burning sequence, of a graph $G=(V,E)$, is a function $f:\{x\in\mathbb{Z}_{+}:x\leq k\}\rightarrow V$ such that Eq. (7) holds. The length of the burning sequence is $k$.
$\bigcup_{i=1}^{k}N_{k-i}[f(i)]=V$ (7)
Definition 9.
Given an input graph, GBP seeks a burning sequence of minimum length.
3 Burning and covering
This section shows how GBP can be reduced to a series of CMCPs.
Definition 10.
The clustered maximum coverage problem (CMCP) receives $k$ clusters as input, each containing subsets of a universe of elements. Its goal is to select one subset from each cluster so that the cardinality of their union is maximized.
Proposition 1.
Maximum coverage problem (MCP) is a particular case of CMCP.
Proof.
MCP receives a collection of subsets and a positive integer $k$ as input. Its goal is to find at most $k$ subsets such that the cardinality of their union is maximized [13]. Notice that an instance for MCP can be transformed into one for CMCP by creating $k$ identical copies of the input collection. Under these conditions, solving CMCP is equivalent to solving MCP. ∎
Since MCP is NP-hard, CMCP is too. The best approximation ratio achievable for the former cannot be greater than $(1-\frac{1}{e})$ unless $P=NP$ [13, 36, 14]. Thus, by Proposition 1, CMCP has the same inapproximability results.
Proposition 2.
GBP can be reduced to a series of CMCPs.
Proof.
Let $V$ be the universe of elements for CMCP, where $G=(V,E)$ is the input graph for GBP. The clusters for the corresponding CMCP are $C_{r}=\{N_{r}[v]:v\in V\},\forall r\in[0,k-1]$, where $k$ is a guess on $b(G)$; breadth-first search [37, 38] can be used to compute the distance between all pairs of vertices efficiently. An optimal solution for CMCP consists of one subset from each cluster. An optimal solution for GBP is obtained by properly concatenating the corresponding vertices of each selected subset (See Definition 8). For instance, the clusters for a path $P_{4}=(\{v_{1},v_{2},v_{3},v_{4}\},\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\}\})$, with $k=b(P_{4})=2$, are $C_{0}=\{N_{0}[v_{1}],N_{0}[v_{2}],N_{0}[v_{3}],N_{0}[v_{4}]\}$ and $C_{1}=\{N_{1}[v_{1}],N_{1}[v_{2}],N_{1}[v_{3}],N_{1}[v_{4}]\}$. An optimal solution for this instance is $\{N_{0}[v_{4}],N_{1}[v_{2}]\}$, which covers all the vertices, and its corresponding optimal burning sequence is $(v_{2},v_{4})$. Naturally, solving CMCP with $k=b(G)$ is necessary for the reduction to be useful. To find that value, we can solve CMCP within a binary search between 1 and $n$ (See Algorithm 1). If $k\geq b(G)$, an optimal solution for CMCP must cover all the vertices (line 8). If the optimal solution for CMCP does not cover all vertices, then $k<b(G)$ (line 10). This way, $k=b(G)$ at some point in the binary search, and the optimal burning sequence is found. ∎
Algorithm 1 GBP as a series of CMCPs
1:A graph $G=(V,E)$
2:An optimal burning sequence
3:$l\leftarrow 1$
4:$h\leftarrow|V|$
5:$d\leftarrow$ Compute the distance matrix
6:while $l\leq h$ do
7:     $k\leftarrow\lfloor(l+h)/2\rfloor$
8:     $s\leftarrow$ ExactSolver$(\text{CMCP}(G,k,d))$
9:     if $s$ burns all vertices then
10:         $h\leftarrow k-1$
11:     else
12:         $l\leftarrow k+1$
13:     end if
14:end while
15:return the best found burning sequence
To further explain this reduction, let us introduce an ILP for GBP as a CMCP (See Expressions (8) to (11)). The clusters are constructed based on the input graph $G=(V,E)$ and a given integer $k$; let us assume that $V=\{v_{1},v_{2},...,v_{n}\}$. This program has $\mathcal{O}(kn)$ binary decision variables and $\mathcal{O}(n)$ constraints. Decision binary variables $x_{rj}$ are used to indicate which subset from each cluster is part of the solution (See Constraints (9)). Actually, this is an ILP for $w$GBP; nevertheless, in this paper, we stick to the original GBP, i.e., $w=1$. Decision binary variables $b_{j}$ indicate which vertices the selected subsets cover. Since the Objective function (8) maximizes the sum over all $b_{j}$, we add Constraints (10) to avoid a vertex $v_{j}$ being marked as covered (i.e., $b_{j}=1$) if it is not an element of some of the selected subsets; notice that |$|\{N_{r}[v]:v\in V\}|=n$ for each cluster $r\in[0,k-1]$. Of course, a burning sequence of length $k$ exists only if $k\geq b(G)$. If so, the burning sequence is codified into variables $x_{ri}$. Explicitly, for $w=1$, the burning sequence is $(v_{g(x_{k-1})},v_{g(x_{k-2})},...,v_{g(x_{0})})$, where $g(x_{r})$ is the index $i$ where $x_{ri}=1$. In case $k<b(G)$, a burning sequence of such length does not exist.
max $\ \ \sum\limits_{v_{j}\in V}b_{j}$ (8)
s.t. $\ \ \sum\limits_{v_{i}\in V}x_{ri}=w$ $\forall r\in[0,k-1]$ (9)
$\ \ b_{j}\leq\sum\limits_{r\in[0,k-1]}\ \ \sum_{v_{i}\in N_{r}[v_{j}]}x_{ri}$ $\forall v_{j}\in V$ (10)
where $\ \ x_{rj}\ ,\ b_{j}\in\{0,1\}$ $\forall r\in[0,k-1],\forall v_{j}\in V$ (11)
The next section introduces a greedy $(1/2)$-approximation algorithm for CMCP and its heuristic adaptation to GBP.
4 Greedy algorithms
This section is divided into two parts. The first introduces a greedy $(1/2)$-approximation algorithm for CMCP. This is used as a heuristic for GBP in the second part.
4.1 A greedy $(1/2)$-approximation algorithm for CMCP
The classical greedy approximation algorithm for MCP iteratively selects the subset that covers as many uncovered elements as possible. This algorithm achieves the best possible ratio of $(1-\frac{1}{e})$ [13, 36, 14]. The greedy approximation algorithm for CMCP follows the same idea. Given $k$ clusters of subsets of a universe of elements, iteratively selects the subset that covers as many uncovered elements as possible among the $k$ clusters (See Algorithm 2). Selecting a subset from a cluster can make it impossible to cover some elements in the following iterations; we refer to these as discarded elements. Figure 2 shows an instance where $OPT/2$ elements are covered at the first iteration and $OPT/2$ are discarded. In fact, $1/2$ is the approximation ratio of this algorithm (See Lemma 1 and Theorem 3), and Figure 2 serves as a tight example.
Algorithm 2 Greedy $(1/2)$-approximation algorithm for CMCP
1:$k$ clusters of subsets, $C_{1},C_{2},...,C_{k}$.
2:A set containing one subset from each cluster.
3:$s,B\leftarrow\emptyset$
4:$I\leftarrow\{1,2,...,k\}$
5:while $I\not=\emptyset$ do
6:     $i\leftarrow\operatorname*{argmax}_{j\in I}\max_{S\in C_{j}}|S-B|$
7:     $I\leftarrow I-\{i\}$
8:     $A\leftarrow\operatorname*{argmax}_{S\in C_{i}}|S-B|$
9:     $s\leftarrow s\cup\{A\}$
10:     $B\leftarrow B\cup A$
11:end while
12:return $s$
Figure 2: In this instance for CMCP, $k=2$, $C_{1}=\{S_{1},S_{2},S_{3}\}$, $C_{2}=\{S_{4}\}$. The optimal solution $\{S_{2},S_{4}\}$ covers $OPT=10$ squares. If the greedy approximation algorithm selects $S_{3}$ at the first iteration, the returned solution will be $\{S_{3},S_{4}\}$, which covers $OPT/2=5$ squares.
Lemma 1.
At each iteration, the number of newly discarded elements cannot exceed the number of newly covered elements.
Proof.
A set $s_{i}$ with maximum coverage $|s_{i}-B|$ is selected at each iteration, where $B$ is the set of covered elements. Thus, the discarded elements, if any, must belong to a subset with coverage smaller than $|s_{i}-B|$. ∎
Theorem 3.
The greedy approximation algorithm for CMCP has an approximation ratio of $1/2$.
Proof.
Let $OPT$ be the size of the optimal solution, let $x_{i}$ be the number of newly covered elements at iteration $i$, let $y_{i}$ be the number of newly discarded elements at iteration $i$, and let $z_{i}$ be the number of available elements (not yet covered nor discarded) at the end of iteration $i$. Since a subset with maximum coverage is selected at each iteration, it must cover at least $1/(k-(i-1))$ of the $z_{i-1}$ available elements; otherwise, $OPT$ would not be optimal (See Eq. (12) and Eq. (13)). By Lemma 1, no more than $x_{i}$ elements can be discarded (See Eq. (14)). Notice that $z_{0}=OPT$ and $z_{k}=0$ because all the available vertices are covered at the last iteration.
$x_{i}\geq\frac{z_{i-1}}{k-(i-1)}$ (12)
$z_{i}=z_{i-1}-x_{i}-y_{i}$ (13)
$y_{i}\leq x_{i}$ (14)
The proof is by contradiction. Let us assume that the algorithm returns a solution that covers less than $OPT/2$ elements, i.e.,
$\sum_{i=1}^{k}x_{i}<\frac{OPT}{2}\ \ .$ (15)
Combining Eq. (14) with Eq. (15), we get
$\sum_{i=1}^{k}y_{i}\leq\sum_{i=1}^{k}x_{i}<\frac{OPT}{2}\ \ .$ (16)
Then, combining Eq. (13) with Eq. (16) we get
$\sum_{i=1}^{k}z_{i}>\sum_{i=1}^{k}z_{i-1}-OPT\ \ .$ (17)
Notice that $\sum_{i=1}^{k}z_{i-1}-\sum_{i=1}^{k}z_{i}=z_{0}-z_{k}$. Since $z_{0}=OPT$ and $z_{k}=0$, Eq. (17) reduces to
$z_{0}-z_{k}=OPT<OPT\ \ ,$ (18)
which is a contradiction. Therefore, $\sum_{i=1}^{k}x_{i}\geq OPT/2$. Namely, the greedy approximation algorithm for CMCP has an approximation ratio of $1/2$. ∎
In the next section, we adapt Algorithm 2 to GBP, resulting in a heuristic algorithm.
4.2 A greedy heuristic for GBP
Algorithm 3 shows the proposed deterministic greedy heuristic (Gr) for GBP, which implements the greedy approximation algorithm for CMCP. In a nutshell, starting from the clusters with larger subsets, i.e., from $k-1$ to 0 (line 3), it iteratively selects the subset that covers as many uncovered elements as possible (line 7). In more detail, Gr receives as input a graph $G=(V,E)$, a guess $k$ on $b(G)$, and the distance matrix $d$, which can be computed independently using breadth-first search. Then, Gr computes the closed neighborhood $N_{k-1}[v]$ of each $v\in V$. In other words, it computes the adjacency list of $G^{k-1}$ (line 2). Afterward, it iteratively selects $k$ subsets that cover as many uncovered vertices as possible in graphs $G^{r}$, where $r$ goes from $k-1$ to 00 (lines 3 to 10). Notice that the adjacency list is updated at each iteration, containing uncovered vertices at a distance at most $r$ from each $v\in V$ (line 5). Besides, every subset $N_{i}[v]$ of each cluster is a superset of each $N_{j}[v]$, where $j<i$. Thus, a subset with maximum coverage among all the clusters is selected at each iteration. Therefore, this is a correct implementation of Algorithm 2. Since the distance matrix $d$ is known, updating the adjacency list can be done in at most $\mathcal{O}(n^{2})$ steps at each iteration. Thus, the overall complexity of Gr is $\mathcal{O}(kn^{2})$; however, if we consider the time required to compute all-pairs’ shortest path, its complexity is $\mathcal{O}(mn+kn^{2})$ as stated in the abstract. By Theorem 3, Proposition 4 follows.
Algorithm 3 Greedy burning (Gr)
1:A graph $G=(V,E)$, its distance matrix $d$, and a guess $k$ on $b(G)$
2:A sequence
3:$s,B\leftarrow\emptyset$
4:$N_{k-1}\leftarrow$ Adjacency list of $G^{k-1}$
5:for $r=k-1$ to 00 do
6:     if $r<k-1$ then
7:         $N_{r}\leftarrow$ Update the adjacency list for $G^{r}$
8:     end if
9:     $v\leftarrow\operatorname*{argmax}_{u\in V}|N_{r}[u]-B|$
10:     $s\leftarrow s\cup\{(k-r,v)\}$
11:     $B\leftarrow B\cup N_{r}[v]$
12:end for
13:return $s$
Proposition 4.
If the subsets Gr selects cover less than $n/2$ vertices, then $k<b(G)$.
Proof.
By Theorem 3, Algorithm 2 returns a solution that covers at least $OPT/2$ elements, where $OPT$ is the optimal solution size. By definition, a burning sequence burns $n$ vertices. So, if $k\geq b(G)$, $OPT=n$, and since Gr correctly implements Algorithm 2, it must return a solution that covers at least $n/2$ vertices. By contrapositive, if Gr returns a solution that covers less than $n/2$ vertices, then $k<b(G)$. ∎
For Gr to be useful, we must know $b(G)$. We can try different guesses $k$ to tackle the issue of not knowing this value in advance. The simpler way is by trying all $k\in[1,n]$. However, a binary search is more efficient. Although Proposition 4 helps find a lower bound for $b(G)$, we didn’t find it useful in practice. Therefore we didn’t integrate it into the binary search (See Algorithm 4), which we now explain. First, the distance between every pair of vertices is computed in $\mathcal{O}(mn+n^{2})$ steps using breadth-first search (line 1). Then, the lower and upper bounds are computed using the burning farthest-first (BFF) approximation algorithm for GBP [12], which runs in $\mathcal{O}(n^{2})$ steps if the computation of all-pairs’ shortest path is not considered (Algorithm 5 shows a detailed pseudocode for BFF using a FIFO queue). Since BFF returns a solution $s_{0}$ (line 2) of length at most $23b(G)-$ 2 [12], the lower bound on $b(G)$ can be safely set to $\lceil(len(s_{0})+2)/3\rceil$ (line 3). Then, Gr is executed with parameters $G$, $k$, and $d$ (lines 6 and 7); GrP will be introduced later, and if an exact solver is used, this is equivalent to Algorithm 1 using tighter bounds. Afterward, every time Gr returns a burning sequence, $h$ is set to $k-1$; otherwise, $l$ is set to $k+1$. This way, Gr is executed up to $\log n$ times with different guesses on $k$ that might approach $b(G)$. Using Gr at line 7, and since $k\leq n$, the complexity of Algorithm 4 is $\mathcal{O}(n^{3}\log n)$.
Algorithm 4 Binary search for GBP
1:A graph $G=(V,E)$
2:A burning sequence
3:$d\leftarrow$ Compute the distance matrix
4:$s_{0}\leftarrow\text{BFF}(G)$
5:$l\leftarrow\lceil(len(s_{0})+2)/3\rceil$
6:$h\leftarrow len(s_{0})$
7:while $l\leq h$ do
8:     $k\leftarrow\lfloor(l+h)/2\rfloor$
9:     $s\leftarrow$ Gr$(G,k,d)$ or GrP$(G,k,d)$ or ExactSolver$(\text{CMCP}(G,k,d))$
10:     if $s$ burns all vertices then
11:         $h\leftarrow k-1$
12:     else
13:         $l\leftarrow k+1$
14:     end if
15:end while
16:return the best found burning sequence
Algorithm 5 Burning Farthest-First (BFF)
1:A graph $G=(V,E)$ and its distance matrix $d$
2:A burning sequence of length at most $23b(G)-$ 2
3:$q\leftarrow$ Empty FIFO queue
4:$v\leftarrow$ Any vertex from $V$
5:$i\leftarrow 1$
6:$s\leftarrow\{(i,v)\}$
7:$q.push(v)$
8:$B\leftarrow\{v\}$
9:for $u\in V$ do
10:     $dist(u)\leftarrow d(u,v)$
11:end for
12:while $B\not=V$ do
13:     $q\_size\leftarrow q.size()$
14:     for $j=1$ to $q\_size$ do
15:         $v\leftarrow q.pop()$
16:         for $u\in N(v)-B$ do
17:              $q.push(u)$
18:              $B\leftarrow B\cup\{u\}$
19:         end for
20:     end for
21:     $v\leftarrow\operatorname*{argmax}_{u\in V}dist(u)$
22:     $i\leftarrow i+1$
23:     $s\leftarrow s\cup\{(i,v)\}$
24:     $q.push(v)$
25:     $B\leftarrow B\cup\{v\}$
26:     for $u\in V$ do
27:         if $d(u,v)<dist(u)$ then
28:              $dist(u)\leftarrow d(u,v)$
29:         end if
30:     end for
31:end while
32:return $s$
Using BFF at line 2 of Algorithm 4 guarantees that a burning sequence of length at most $23b(G)-$ 2 is found [12]. Nevertheless, in the worst scenario, Gr might not find a better solution than that given by BFF. In fact, it might not return a burning sequence for paths and cycles (See Proposition 5 and Corollary 1). Since there are exact $\mathcal{O}(n)$ algorithms for GBP over these graphs’ families [15], the behavior of Gr over them might hint at its theoretical limitations.
Proposition 5.
Gr might not return a burning sequence for a path $P$.
Proof.
It is well known that $b(P_{n})=\lceil n^{1/2}\rceil$ for any path $P_{n}$ of order $n$ [1]. So, given a path $P_{n}$, we can run Gr with $k=b(P_{n})$, which is equivalent to execute Algorithm 4 with $l=h=b(P_{n})$. For the rest of the proof, let us consider paths $P_{n}$ with the largest possible order for any given burning number, namely, paths $P_{n}$ with $n=z^{2}$ for any $z\in\mathbb{Z}^{+}$. Also, observe that the maximum number of vertices a vertex $v$ can cover in a path is $|N_{r}[v]|=2r+1$, for any given $r$. The proof is divided into three parts. The first considers the initial iteration of the for loop of Gr ($r=k-1$); the second considers iterations $r=k-2$ to $r=2$, and the last one considers iterations $r=1$ and $r=0$. In the first iteration, $r=k-1$, Gr selects a vertex that covers exactly $12r+$ 1 vertices because the path’s length, $(r+1)^{2}$, is bounded by $12r+$ 1, i.e.,
$|P_{n}|=(r+1)^{2}\geq 2r+1\ \ .$ (19)
At iterations $r=k-2$ to $r=2$, Gr selects a vertex with maximum coverage from the largest subpath induced by removing the covered vertices. The largest subpath has at least $12r+$ 1 vertices, i.e.,
$\max\{|P_{a}|,|P_{b}|\}\geq\Bigl{\lceil}\frac{(r+2)^{2}-(2(r+1)+1)}{2}\Bigr{\rceil}\geq 2r+1\ \ ,$ (20)
where $(r+2)^{2}$ is the number of uncovered vertices two iterations before, and $12(r+1)+$ 1 is the number of newly covered vertices at the previous iteration. Thus, the numerator is the number of uncovered vertices at the beginning of iteration $r$. This way, at iteration $r$, Gr selects a vertex that covers $12r+$ 1 vertices, none of which were covered in the previous iterations. Notice that ties can be broken at random. By mathematical induction, Eq. (20) holds for any integer $r\geq 2$. At iteration $r=1$, the number of previously covered vertices equals $\sum_{r=2}^{k-1}(2r+1)=k^{2}-4$. Namely, at this point, only four vertices are left uncovered. By direct inspection, there will be cases where only two or three new vertices can be covered at iterations $r=1$ and $r=0$, leaving the rest uncovered (See Figure 3). Therefore, the constructed solution might not be a burning sequence.
Figure 3: A non-burning sequence returned by Gr over $P_{25}$ with $k=b(P_{25})=5$. At the end, two vertices are uncovered.
Corollary 1.
Gr might not return a burning sequence for a cycle $C$.
Proof.
Since a cycle $C_{n}$ is a path $P_{n}$ with its endpoints joined, it follows that $b(C_{n})=\lceil n^{1/2}\rceil$ [1]. So, we can run Gr with $k=b(C_{n})$, which is equivalent to execute Algorithm 4 with $l=h=b(C_{n})$. The first vertex selected by Gr maximizes the coverage. Then, removing the covered vertices induces a path $P$. By Proposition 5, Gr might return a non-burning sequence of length $b(C_{n})-1$ over $P$, which is concatenated to the first selected vertex, becoming a non-burning sequence of length $b(C_{n})$ for $C_{n}$. ∎
By Proposition 5 and Corollary 1, Gr might not seem very appealing. However, as reported in Section 5, it performed unexpectedly well over benchmark and synthetic graphs. Anyhow, we can do better by repeating Gr up to $n$ times and selecting a different initial vertex every time; we refer to this heuristic as GrP (Algorithm 6). Once the first burning sequence of length $k$ is found (line 3), exploring the remaining vertices is unnecessary. Nevertheless, the complexity of GrP embedded into a binary search (Algorithm 4) is $\mathcal{O}(n^{4}\log n)$, which is impractical for sufficiently large graphs.
Algorithm 6 Greedy burning plus (GrP)
1:A graph $G=(V,E)$, its distance matrix $d$, and a guess $k$ on $b(G)$
2:A sequence
3:for $v\in V$ do
4:     $s\leftarrow Gr(G,k,d)$ setting $(1,v)$ into the sequence
5:     if $s$ burns all vertices then
6:         return $s$
7:     end if
8:end for
9:return the best found sequence
5 Experimental results
This section reports the execution time and the size of the burning sequences returned by Algorithm 4 using Gr (Algorithm 3), GrP (Algorithm 6), and off-the-shelf optimization software (implementing Expressions (8) to (11)) over benchmark and synthetic graphs. All the real-world networks were obtained from the network repository [39] and the Stanford large network dataset collection [40]. Gr and GrP were implemented in C++ and compiled with GNU GCC. The ILP from Section 3 (Expressions (8) to (11)) was solved to optimality by Gurobi 11.0.0 [41] using default parameters and an aggressive presolve method. The execution times, even on small graphs, reported by Gurobi using the ILP from Section 2 (Expressions (1) to (6)) were very unpractical; thus, we did not report them. Table 1 shows all experiments’ hardware and software specifications. The implemented code can be consulted at https://github.com/jesgadiaz/GreedyBurning.
Table 1: Hardware and software setup
For all experiments
    CPU Intel Core i5-10300H
    CPU frequency 2.5GHz
    CPU cores 4
    RAM 32GB
    OS Windows 11
For Gr and GrP
    Programming language C++
    Compiler GNU GCC
For ILP
    Solver Gurobi 11.0.0
    Presolve 2
    MIPGap 0%
    Heuristics 5%
Table 2 (small real-world graphs), Table 3 (medium and large real-world graphs), and Table 4 (square grids) show the execution time and the length of the solutions found by Gr, GrP, and Gurobi within a binary search (Algorithm 4). Column BFF reports the bounds used in the binary search; we used BFF algorithm to set $l=\lceil(len(s_{0})+2)/3\rceil$ and $h=len(s_{0})$, where $len(s_{0})$ is the length of the burning sequence returned by this algorithm; for reproducibility purposes, the first vertex selected by BFF is the one with the smallest label. Column BFS reports the time required to compute the distance matrix using breadth-first search (all reported times are in seconds). Using the ILP from Section 3 (Expressions (8) to (11)), Gurobi corroborated all the previously reported optimal solutions and found six new ones (tvshow, government, crocodile, DD6, DD349, and DD687). According to these results, the introduced ILP seems advantageous compared to the state-of-the-art ILP and CSPs [35]. For instance, Gurobi found the optimal burning sequence of tvshow in 473 seconds ($\sim$8 minutes) using the ILP from Section 3, while the same software could not find it in more than 48 hours using the state-of-the-art ILP and CSPs [35]. Besides, while the largest instance solved using the state-of-the-art formulations has 5908 vertices, Gurobi solved an instance of order 11631 in 1989 seconds ($\sim$33 minutes) using our formulation; we get an out-of-memory error for larger graphs. As expected, Gr was the fastest among all the tested algorithms. Its larger execution time was 1587 seconds ($\sim$26.5 minutes) over a graph of order 102400. A dash at Tables 2 to 4 indicates that the corresponding procedure was not executed because the memory requirements exceeded the hardware’s capacity or a time limit of 10 hours was estimated to be surpassed.
Table 2: Experimental results over small real-world networks. This table reports the execution time and quality of solutions found by Gr (Algorithm 3), GrP (Algorithm 6), and Gurobi (using Expressions (8) to (11)), all embedded into a binary search (Algorithm 4). Column BFF reports the time consumed by BFF algorithm and the bounds used in the binary search. Column BFS reports the time consumed by BFS to compute the distance matrix. Optimal solutions are emphasized, and all times are in seconds; a time of 0.000 represents a time below milliseconds.
Graph BFF BFS Gurobi Gr GrP
name $n$ $m$ $b(G)$ $l$ $h$ $t$ $t$ $t$ size $t$ size $t$
karate 34 78 3 2 4 0.000 0.000 0.004 3 0.000 3 0.000
chesapeake 39 170 3 1 3 0.000 0.000 0.009 3 0.000 3 0.000
dolphins 62 159 4 2 6 0.000 0.000 0.036 4 0.000 4 0.010
rt-retweet 96 117 5 2 6 0.000 0.000 0.035 5 0.000 5 0.000
polbooks 105 441 4 2 5 0.000 0.000 0.047 4 0.000 4 0.015
adjnoun 112 425 4 2 5 0.000 0.000 0.044 4 0.004 4 0.015
ia-infect-hyper 113 2196 3 1 3 0.000 0.015 0.04 3 0.004 3 0.016
C125-9 125 6963 3 1 3 0.000 0.031 0.094 3 0.000 3 0.063
ia-enron-only 143 623 4 2 5 0.000 0.000 0.083 4 0.000 4 0.046
c-fat200-1 200 1534 7 3 7 0.000 0.015 0.235 7 0.000 7 0.220
c-fat200-2 200 3235 5 2 5 0.000 0.031 0.177 5 0.000 5 0.220
c-fat200-5 200 8473 3 1 3 0.000 0.062 0.079 3 0.004 3 0.095
sphere 258 1026 7 3 9 0.000 0.015 2.0 7 0.000 7 0.282
DD244 291 822 7 4 11 0.000 0.015 0.739 7 0.000 7 0.727
ca-netscience 379 914 6 3 8 0.000 0.031 0.632 7 0.003 6 0.644
infect-dublin 410 2765 5 2 6 0.000 0.062 0.614 5 0.015 5 1.2
c-fat500-1 500 4459 9 4 11 0.000 0.106 1.2 9 0.015 9 2.5
c-fat500-2 500 9139 7 3 8 0.000 0.203 1.0 7 0.015 7 3.0
c-fat500-5 500 23191 5 2 5 0.000 0.456 0.868 5 0.015 5 3.6
bio-diseasome 516 1188 7 5 13 0.000 0.046 1.6 7 0.031 7 2.0
web-polblogs 643 2280 5 3 8 0.000 0.109 2.2 6 0.046 5 6.0
DD687 725 2600 7 4 10 0.000 0.125 12 8 0.062 8 8.0
rt-twitter-copen 761 1029 7 3 9 0.000 0.093 4.7 7 0.078 7 9.3
DD68 775 2093 9 5 14 0.000 0.124 10 10 0.051 9 16
ia-crime-moreno 829 1475 7 3 8 0.000 0.112 12 7 0.078 7 20
DD199 841 1902 12 6 16 0.000 0.125 12 13 0.062 12 7.0
soc-wiki-Vote 889 2914 6 3 8 0.000 0.171 7.7 6 0.094 6 16
DD349 897 2087 12 6 18 0.000 0.157 12 13 0.062 12 20
DD497 903 2453 10 6 16 0.000 0.160 15 12 0.059 11 23
socfb-Reed98 962 18812 4 2 5 0.000 0.735 7.3 4 0.064 4 16
lattice3D 1000 2700 10 4 12 0.000 0.203 218 10 0.094 10 38
bal-bin-tree-9 1023 1022 10 4 10 0.000 0.125 3.1 10 0.031 10 5.8
delaunay-n10 1024 3056 9 4 11 0.000 0.236 33 10 0.078 9 27
stufe 1036 1868 12 5 15 0.000 0.176 122 12 0.078 12 22
lattice2D 1089 2112 13 7 19 0.000 0.187 172 14 0.078 13 40
bal-ter-tree-6 1093 1092 7 3 7 0.000 0.157 2.2 7 0.015 7 3.8
email-univ 1133 5451 5 2 6 0.000 0.359 13 5 0.094 5 23
econ-mahindas 1258 7513 5 2 6 0.000 0.488 54 5 0.109 5 31
ia-fb-messages 1266 6451 5 2 6 0.000 0.461 17 5 0.141 5 42
bio-yeast 1458 1948 9 4 11 0.000 0.343 29 9 0.234 9 143
Table 3: Experimental results over medium and large real-world networks. This table reports the execution time and quality of solutions found by Gr (Algorithm 3), GrP (Algorithm 6), and Gurobi (using Expressions (8) to (11)), all embedded into a binary search (Algorithm 4). Column BFF reports the time consumed by BFF algorithm and the bounds used in the binary search. Column BFS reports the time consumed by BFS to compute the distance matrix. Optimal solutions are emphasized, and all times are in seconds; a time of 0.000 represents a time below milliseconds.
Graph BFF BFS Gurobi Gr GrP
name $n$ $m$ $b(G)$ $l$ $h$ $t$ $t$ $t$ size $t$ size $t$
tech-routers-rf 2113 6632 6 3 8 0.000 1.0 89 6 0.508 6 193
chameleon 2277 31421 6 3 8 0.000 3.4 61 6 0.659 6 350
tvshow 3892 17262 9 5 13 0.000 4.0 473 10 1.9 9 5201
facebook 4039 88234 4 2 5 0.000 14 29 4 0.533 4 369
DD6 4152 10320 16 9 25 0.000 3.4 1047 17 1.4 17 2868
squirrel 5201 198493 6 2 6 0.000 38 499 6 3.2 6 7394
politician 5908 41729 7 3 9 0.000 13 1239 7 3.7 7 8935
government 7057 89455 6 3 7 0.000 28 1233 6 4.2 6 10299
crocodile 11631 170918 6 3 7 0.000 84 1989 6 12 6 46798
athletes 13866 86858 ? 3 8 0.000 65 - 7 25 - -
company 14113 52310 ? 4 10 0.015 50 - 9 26 - -
musae-facebook 22470 171002 ? 4 11 0.015 208 - 8 72 - -
new-sites 27917 206259 ? 3 9 0.031 322 - 8 109 - -
deezer-europe 28281 92752 ? 4 12 0.031 208 - 10 105 - -
gemsec-deezer(RO) 41773 125826 ? 4 12 0.016 463 - 10 266 - -
gemsec-deezer(HU) 47538 222887 ? 4 10 0.046 780 - 8 376 - -
artist 50515 819306 ? 3 8 0.031 2035 - 6 418 - -
gemsec-deezer(HR) 54573 498202 ? 3 8 0.047 1631 - 7 426 - -
Table 4: Experimental results over square grids. This table reports the execution time and quality of solutions found by Gr (Algorithm 3), GrP (Algorithm 6), and Gurobi (using Expressions (8) to (11)), all embedded into a binary search (Algorithm 4). Column BFF reports the time consumed by BFF algorithm and the bounds used in the binary search. Column BFS reports the time consumed by BFS to compute the distance matrix. Optimal solutions are emphasized, and all times are in seconds; a time of 0.000 represents a time below milliseconds.
Square grid BFF BFS Gurobi Gr GrP
name $n$ $m$ $b(G)$ $l$ $h$ $t$ $t$ $t$ size $t$ size $t$
10x10 100 180 6 3 8 0.000 0.001 0.161 7 0.015 6 0.031
20x20 400 760 10 5 13 0.000 0.027 5.2 11 0.017 10 0.897
30x30 900 1740 12 6 17 0.000 0.14 177 13 0.078 12 13
40x40 1600 3120 15 7 20 0.000 0.45 23742 16 0.191 15 92
50x50 2500 4900 ? 8 23 0.000 1.1 - 18 0.484 17 439
60x60 3600 7080 ? 9 26 0.011 2.2 - 21 1.1 19 1202
70x70 4900 9660 ? 10 29 0.000 4 - 22 1.5 21 3020
80x80 6400 12640 ? 11 31 0.000 7 - 25 3.2 24 5688
90x90 8100 16020 ? 12 34 0.000 11 - 28 5.1 25 12492
100x100 10000 19800 ? 12 36 0.013 17 - 28 5.8 27 32606
110x110 12100 23980 ? 13 38 0.013 25 - 30 10 - -
120x120 14400 28560 ? 14 41 0.015 36 - 32 15 - -
130x130 16900 33540 ? 15 43 0.015 50 - 34 20 - -
140x140 19600 38920 ? 16 46 0.046 69 - 35 27 - -
150x150 22500 44700 ? 16 48 0.046 93 - 37 35 - -
160x160 25600 50880 ? 17 51 0.062 123 - 39 44 - -
170x170 28900 57460 ? 18 53 0.063 161 - 41 56 - -
180x180 32400 64440 ? 19 55 0.063 208 - 42 69 - -
190x190 36100 71820 ? 19 57 0.094 259 - 44 85 - -
200x200 40000 79600 ? 20 60 0.141 322 - 45 123 - -
210x210 44100 87780 ? 20 60 0.126 397 - 47 124 - -
220x220 48400 96360 ? 21 63 0.157 487 - 48 150 - -
230x230 52900 105340 ? 22 64 0.181 591 - 50 215 - -
240x240 57600 114720 ? 23 67 0.188 713 - 51 211 - -
250x250 62500 124500 ? 23 67 0.345 838 - 52 300 - -
260x260 67600 134680 ? 23 69 0.219 1019 - 54 351 - -
270x270 72900 145260 ? 24 70 0.464 1174 - 55 402 - -
280x280 78400 156240 ? 25 73 0.456 1388 - 57 471 - -
290x290 84100 167620 ? 25 73 5.2 1627 - 58 641 - -
300x300 90000 179400 ? 25 74 19 1954 - 59 983 - -
310x310 96100 191580 ? 26 76 53 2306 - 60 1522 - -
320x320 102400 204160 ? 26 78 35 2687 - 62 1587 - -
6 Discussion
GBP is an NP-hard minimization problem with application to information spreading problems. Some heuristics and metaheuristics have been proposed to find feasible solutions. However, research about more straightforward greedy approaches wasn’t fully addressed. This paper introduces a simple deterministic greedy heuristic, Gr (Algorithm 3), tailored for the CMCP, a problem to which GBP is reduced if $b(G)$ is known in advance; the issue of not knowing this value is tackled by using binary search (Algorithm 4). Although Gr cannot guarantee to find optimal solutions for paths and cycles for which exact $\mathcal{O}(n)$ algorithms exist, it performed unexpectedly well over real-world and synthetic graphs. Namely, for a benchmark dataset of small real-world networks, Gr found 31 optimal solutions out of 40, each in no more than 0.234 seconds. Its plus version, GrP, found 38 out of 40 optimal solutions, each in less than 2.4 minutes (See Table 2). Regarding larger real-world networks, Gr found 7 out of 9 optimal solutions and nine new best-known solutions, each in less than 426 seconds ($\sim$7 minutes); GrP found 8 out of 9 optimal solutions, though it required up to 46798 seconds ($\sim$13 hours) (See Table 3). Regarding synthetic graphs, square grids seem challenging to solve because they have a relatively large burning number. Over these instances, although Gr could not find optimal solutions, GrP could. Nevertheless, the maximum execution time for Gr was 1587 seconds ($\sim$26.5 minutes) over a grid of order 102400, while GrP required 32606 seconds ($\sim$9.1 hours) over a grid of order 10000. Gurobi became unpractical on grids of order at least 1600 (See Table 4). By using Gurobi, Algorithm 4 corroborated all the previously known optimal solutions and found six new ones (tvshow, government, crocodile, DD6, DD349, and DD687); the implemented ILP (Expressions (8) to (11)) showed a better performance than the state-of-the-art mathematical formulations for GBP (See Tables 2 to 4).
7 Conclusions
This paper shows how GBP can be reduced to a series of CMCPs (See Proposition 2, Algorithm 1, and Expressions (8) to (11)). Through experimental results, we show that a deterministic greedy $(1/2)$-approximation algorithm for CMCP (Algorithm 2) can be used to find feasible solutions for GBP. Although the deterministic greedy heuristic for GBP, Gr (Algorithm 3), might not find optimal solutions of paths and cycles, it might have interesting theoretical properties over other graph families; we leave that research for future work. Nevertheless, the proposed heuristic performed unexpectedly well over real-world and synthetic graphs, finding most of the optimal and best-known solutions in reasonable amounts of time.